# Benchmarking Multi-Agent Preference-based Reinforcement Learning for Human-AI Teaming

### v1

@@ 5.1 Human-AI PbRL Cooperation Game
  • $T(\cdot|\cdot,\cdot,\cdot)$ a conditional distribution on the next world state, given previous state and action for both agents, i.e.,
    $T(s^{\prime}|s,a^{H},a^{R})$;
@@
  • $R_{\theta}(\cdot,\cdot,\cdot)$ a reward function that maps world states, joint actions to reward values according to H’s preference on the team objective, H’s objectives, R’s objectives and preference over R’s behavior;
  • $\gamma$ a discount factor : $\gamma\in[0,1]$
@@
The reward function $R$ is set by the human agent defining the team objectives and human-agent acceptable policy pairs $(\pi_{H},\pi_{R})$. The game begins at the initial state $s_{o}\in S$. The human agent pre-computes acceptable solutions to the game with assumed knowledge on AI partner’s capabilities. Suppose the computed solution set is $\Pi=(\pi_{H},\pi_{R})_{0},\pi_{H},\pi_{R})_{1}\cdots\pi_{H},\pi_{R})_{n}$. For simplicity, suppose the solution set is singleton. The human takes an action following their policy $\pi_{H}$. The robot takes a random action in $\mathcal{A}^{R}$. As the game proceeds the robot builds a bank of trajectories from past experiences. Following PbRL setup, the agent samples two trajectories from this bank and queries the human agent for their feedback. The human agent uses their internal reward function $R$ to provide their preference label. A behavior pair where the shown human actions align with the true human actions from $\pi_{H}$ and the shown robot actions align with human expected $\pi_{R}$ such that resultant return of the trajectory is higher than the alternate is preferred.
While the policies can be arbitrary functions of history to next action as $\pi_{H}:[\mathcal{A}^{H}\times\mathcal{A}^{R}\times S]*\rightarrow\mathcal{A}^{H}$ (and similarly for robot R), we restrict the policy to be defined on states (thereby assuming the dynamics, reward function and the optimal policy pair is Markovian with respect to the world state).
5.2 Flexibility of the Human-in-the-Loop
@@
It is important to note that for our human-AI teaming PbRL setup, we only consider the case where the AI agent’s access to the underlying reward function is restricted. That is, the agent does not experience a sampled reward from the environment, and instead, depends on an explicit reward learning procedure (through human feedback) to realize its objectives in the team. Assume the set $\Pi=\{\Pi_{1},\Pi_{2},\cdots\Pi_{n}\}$ where $\Pi_{i}=(\pi_{i}^{H},\pi_{i}^{R})$ be the possible solutions of the Cooperation Game defined in Section 5.1 as can be computed by the human. We define human flexibility over their policy, i.e. $\{\pi_{i}^{H}\}_{\{i=1\cdots n\}}$ as :
Definition 5.2.
@@
Given $\Pi$ solution set, the human in the loop is said to be $\epsilon$-H-flexible if
|$\left|\{\pi_{k}^{H}\,|\,k=1\,\text{to}\,n\}\right|=\epsilon-1$
@@
The definition of H-flexible (5.2) human in the loop helps us recognize the rigidity of the human agent over a strategy for the two agents. A large $\epsilon\rightarrow|\Pi|$ implies that the human is willing to execute any one of the policies in $\Pi_{H}$ subsequently providing the robot to converge to a larger solution space (policies $\pi_{k}^{R}$). The nature and requirements of the cooperation for the given team-goal can be helpful in determining the values of $\epsilon$. For instance, $\epsilon=0$ is useful when considering human cognitive limitations, thereby the human is able to come up with a single policy for themselves.
@@
When $\epsilon>1$ then the human agent has a choice to make in selecting which policy $\pi_{k}^{H}$ they will follow. In this work, we assume that the AI agent’s objective is to be conformant to all the human policies. This choice is motivated from how prior works in Human-AI cooperation have attempted to learn the robot’s policy [53]. Other alternatives are: the agent is compatible to at least one of the human policies, a subset of the human policies etc. exist which would also impact the training regime of the RL agent (i.e. which policy is used by the human agent when the Robot is learning its policy?). Following prior works [53, 54] in learning a response policy, during training we randomly sample the human’s policy which is used for an entire episode.
5.3 Access to Human Policy
@@
Based on the Human-AI Team PbRL cooperative game describe in Section 5.1, the human and the robot take actions in a shared environment. This implies that for the duration of training, in addition to the preference feedback, the framework assumes the human to sample their optimal action $a_{h}$ for a world state $s$. That is, the agent gains indirect access to the human policy by virtue of the shared world state and the fact that human-AI coexist in the shared world. While this is the default operating assumption through this work, it has higher requirements from the human agent as compared to single-agent PbRL. For example, in single-agent PbRL the human is required to be present in the loop during agent’s training procedure and provide feedback, whereas in the current setup, in addition to the preference feedback the human is also required to follow their policy throughout the training process (which can be millions of steps).
@@
We can reinterpret the Human-AI Team PbRL cooperative game as follows: the agent has access to an environment simulator which can simulate the agent’s actions and the human’s actions. Now, the agent uses the policy $\pi_{R}$ being learned and an imagined human policy $\pi_{R}^{h}$ to generate team behaviors. These behaviors can be stored in a trajectory buffer and queried to the human actor. The advantage of this re-interpretation is that while the human agent is present in the loop they are not required to execute millions of actions during agent’s training. The disadvantage is, the agent has to model the human’s policy in addition to the human’s preferences. This re-interpretation of the cooperative game can enable us to study the performance of the PbRL agent when it has budgeted access to human actions. Full access is equivalent to the default interpretation, 50% access is when the agent can query for the human’s actions for the first half of the training process and 0% access implies the agent cannot query human’s actions at all. In this work, the agent defaults to using random actions as imagined human actions when it exhausts the human-action query budget. Techniques like behavior-cloning [6] can be leveraged to independently model the human’s policy along with the PbRL step for modeling human preferences in future work.
5.4 Specified Orchestration
@@
Recall our Two Arm Peg-In-Hole example in Section 2. It is clear that since the human agent defines the team objective through reward function $R$, they are restricting the possible robot behaviors to achieve the team goal. For example, the robot arm should approach very slowly towards the human controlled arm because of human’s preference for safety. It becomes extremely challenging if the agent does not have access to the human’s policy. Further, as described in Section 5.2, if the human is switching their policy in the feasible team behavior set during the agent’s training process, the agent now has an additional objective of learning a conformant response policy that is compatible with all of the policies in $\epsilon$-Human Flexible solution set. Our empirical results in sections 6 show that the agent’s performance in policy learning is significantly diminished in the absence of these assumptions.
Definition 5.3.
@@
Specified Orchestration is a special case of the Human-AI Team Cooperative game $M=(S,\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$ where the human agent (H) is 1-H-flexible with the team behavior solution $(\pi_{H},\pi_{R})$ and $\pi_{H}$ is accessible by robot R.
@@
Note that this is a special case of the cooperative game where we have specific assumptions on the human’s preference on the team behavior and access to human’s actions. While these assumptions are quite expensive to meet, they play an important role in benchmarking Human-AI Team PbRL algorithms. Specified Orchestration provides a loose upper-bound performance by assuming away other challenging aspects of Human-AI interaction except preference elicitation, reward learning and subsequent policy learning.
6 Benchmarking Human-AI Team PbRL
In this section we will discuss the proposed benchmark suite (Section 6.1) and our experiments showcasing performance of Human-AI Team PbRL across the two critical aspects (as discussed in Sections 6.4 and 6.3) i.e. impact of human flexibility on team behavior and the impact of providing access to human policy to the PbRL agent.
@@ 6 Benchmarking Human-AI Team PbRL
Further, our discussion will highlight the challenges associated with using PbRL for Human-AI team cooperative game and empirically answer the following questions :
@@
    What is the relative performance when the human partner can choose from multiple possible policies to follow, for accomplishing the team task?
@@
    In Human-AI team PbRL, what is the performance of the adapted single-agent PbRL algorithms with maximal assumptions on access to human policy and $\epsilon=1$-H-flexible human (Specified Orchestration)?
@@
    How do PbRL algorithms perform with a budgeted access and no access to human’s policy?
Question (1) supports our discussion on Human Flexibility, (2) highlights the case where PbRL algorithms are effective making Specified Orchestration a loose upper bound performance for studying Human-AI team PbRL performance. Question (3) discusses the special case of limiting agent’s access to human policy thereby highlighting the need for an explicit modeling of human policy.
In the subsequent sections we first study (Q2) which provides us with an upper bound performance of PbRL algorithms followed by (Q3) where we find that the adapted PbRL algorithms fail without access to human policy. Finally, we study (Q1) under full-access to human policy with a $\epsilon>1$-H-flexible solution set.
@@
We refer the readers to the supplementary material (Sections A, B, and C) for detailed discussion on algorithms, the human policy and preference oracles, domain construction and the respective hyperparameters used for the experiments.
Figure 2: MA Highway domain - human and the AI agent cars are shown one behind the other.
Figure 3: MA MuJoCo domains - (from L to R) Cheetah, Ant, Walker, Swimmer, & Hopper. The action space for the joints in each of the domains has been split between the human and the AI agent.

### v2

@@ 4 PbRL for Human-AI Teaming
  • $T(\cdot|\cdot,\cdot,\cdot)$ a conditional distribution on the next world state, given previous state and action for both agents, i.e., $T(s^{\prime}|s,a^{H},a^{R})$;
@@
  • $R_{\theta}(\cdot,\cdot,\cdot)$ a reward function that maps world states, joint actions to reward values according to H’s preference on the team objective, H’s objectives, R’s objectives and team behavior;
  • $\gamma$ a discount factor : $\gamma\in[0,1]$;
  • $\mathcal{H}_{\Pi}$ Human Flexibility: $|\Pi_{H}|$, i.e., set of policies admissible to the human (see Definition 4.2)
  • $\mathcal{H}_{\mathcal{B}}$ Access to Human Policy: $\mathcal{H}_{\mathcal{B}}\in[0,1]$, i.e., R’s budget for querying the human policy (see Definition 4.3)
@@
The reward function $R$ is set by the human agent defining the team objectives and human-agent acceptable policy pairs $(\pi_{H},\pi_{R})$. Note that in teaming settings, $R$ can be decomposed further into feasibility or environment constraints, task reward, and human preferences. However, due to their inherent limitation, PbRL frameworks have only looked at jointly learning a single reward function, and it is also non-trivial to combine multiple rewards otherwise. Now, consider that the game begins at the initial state $s_{o}\in S$. The human agent pre-computes acceptable solutions to the game with assumed knowledge on AI partner’s capabilities. Suppose the computed solution set is $\Pi=(\pi_{H},\pi_{R})_{0},(\pi_{H},\pi_{R})_{1}\cdots(\pi_{H},\pi_{R})_{n}$. For simplicity, suppose the solution set is singleton, i.e. $|\pi_{H}|=1$, and the agent has complete access to human policy, i.e., $\mathcal{H}_{\mathcal{B}}=1$. The human takes an action following their policy $\pi_{H}$. The robot takes a random action in $\mathcal{A}^{R}$. As the game proceeds, robot builds a trajectory bank from past experiences. Following PbRL setup, the agent samples two trajectories from this bank and queries for human feedback. The human agent uses their internal reward function $R$ to provide their preference label. A behavior pair where the shown human actions align with the true human actions from $\pi_{H}$ and the shown robot actions align with human expected $\pi_{R}$ such that resultant return of the trajectory is higher than the alternate is preferred. While policies can be arbitrary functions of history to next action as $\pi_{H}:[\mathcal{A}^{H}\times\mathcal{A}^{R}\times S]*\rightarrow\mathcal{A}^{H}$ (and similarly for robot R), we restrict the policy to be defined on states (thereby assuming the dynamics, reward function and the optimal policy pair is Markovian w.r.t. the world state).
Flexibility of the Human-in-the-Loop
@@
It is important to note that for our human-AI teaming PbRL setup, we only consider the case where the AI agent’s access to the underlying reward function is restricted. That is, the agent does not experience a sampled reward from the environment, and instead, depends on an explicit reward learning procedure through human feedback. Assume the set $\Pi=\{\Pi_{1},\Pi_{2},\cdots\Pi_{n}\}$ where $\Pi_{i}=(\pi_{H},\pi_{R})_{i}$ be the possible solutions of the Cooperation Game defined in Section 4 as can be computed by the human. We define Human Flexibility over their policy, i.e. $\mathcal{H}_{\Pi}$ as :
Definition 4.2
@@
Given the complete solution set $\Pi$ , the human-in-the-loop is said to be $\epsilon$-H-flexible if
$\mathcal{H}_{\Pi}=\left|\{{\pi_{H}}_{k}\,|\,k=1\,\text{to}\,n\}\right|=\epsilon-1$
@@
The definition of H-flexible (4.2) human in the loop helps us recognize the rigidity of the human agent over a strategy for the two agents. A large $\epsilon\rightarrow|\Pi_{H}|$ implies that the human is willing to execute any one of the policies in $\Pi_{H}$ subsequently providing the robot to converge to a larger solution space (policies ${\pi_{R}}_{k}$). The nature and requirements of the cooperation for the given team-goal can be helpful in determining the values of $\epsilon$. For instance, $\epsilon=0$ is useful when considering human cognitive limitations, thereby the human is able to come up with a single policy for themselves.
@@
When $\epsilon>1$, the human agent can select which policy ${\pi_{H}}_{k}$ they will follow. In this work, we assume that the AI agent’s objective is to be conformant to all human policies. This choice is motivated from how prior works in Human-AI cooperation have attempted to learn the robot’s policy (Strouse et al. 2021). Other alternatives are: the agent is compatible to at least one of the human policies, a subset of the human policies etc. exist which would also impact the training regime of the RL agent (i.e. which policy is used by the human agent when the robot is learning its policy).
Access to Human Policy
@@
Based on the Human-AI Team PbRL Cooperation Game describe in Section 4, the human and the robot take actions in a shared environment. This implies that for the duration of training, in addition to the preference feedback, the framework assumes the human to sample their optimal action $a_{h}$ for a world state $s$. That is, the agent gains indirect access to the human policy by virtue of the shared world state and the fact that both agents coexist in the shared world. While this is the default operating assumption through this work, it has higher requirements from the human agent as compared to single-agent PbRL. For example, in single-agent PbRL the human is required to be present in the loop during agent’s training procedure and provide feedback, whereas in the current setup, in addition to the preference feedback, the human is also required to follow their policy throughout the training process (which can be millions of steps).
@@
We can re-interpret the Human-AI Team PbRL Cooperation Game as follows: the agent has access to an environment simulator which can simulate the agent’s actions and the human’s actions. Now, the agent uses the policy $\pi_{R}$ being learned and an imagined human policy $\pi_{R}^{h}$ to generate team behaviors. These behaviors can be stored in a trajectory buffer and queried to the human actor. The advantage of this re-interpretation is that while the human agent is present in the loop, they are not required to execute millions of actions during agent’s training. The disadvantage is, the agent has to model the human’s policy in addition to the human’s preferences. This re-interpretation of the game can enable us to study the performance of the PbRL agent when it has budgeted access to human actions. Formally,
Definition 4.3
Given, the agent’s total training budget of $N$ episodes, and the budget to query human policy for $\mathcal{Q}$ episodes, agent’s access to human policy, $\mathcal{H}_{\mathcal{B}}$, is defined as:
$\mathcal{H}_{\mathcal{B}}=\mathcal{Q}\big{/}N,$
Full access is equivalent to the default interpretation, 50% access is when the agent can query for the human’s actions for the first half of the training process and 0% access implies the agent cannot query human’s actions at all. In this work, the agent defaults to using random actions as imagined human actions when it exhausts the human-action query budget. Techniques like Behavior Cloning (Carroll et al. 2019) can also be leveraged to independently model the human’s policy along with the PbRL step in future work.
Specified Orchestration
@@
It is clear that since the human agent defines the team objective through reward function $R$, they are restricting the possible robot behaviors to achieve the team goal. It becomes extremely challenging to learn this reward function if the agent does not have access to the human’s policy. Further, as described in Section 4, if the human is switching their policy in the feasible team behavior set during the agent’s training process, the agent now has an additional objective of learning a conformant response policy that is compatible with all of the policies in $\epsilon$-Human Flexible solution set. Empirical results in Section 6 show that the agent’s performance in policy learning is significantly diminished in the absence of these assumptions. Hence, we formally call this constrained case Specified Orchestration, and it can be defined as:
Definition 4.4
@@
Specified Orchestration is a special case of the Human-AI Team Cooperation Game $M=(S,\{{\mathcal{A}^{H},\mathcal{A}^{R}}\},T,R_{\theta},\gamma,\mathcal{H}_{\Pi},\mathcal{H}_{\mathcal{B}})$ where the human agent (H) is 1-H-flexible with the team behavior solution $(\pi_{H},\pi_{R})$ and $\pi_{H}$ is completely accessible by robot R, i.e., $|\Pi_{H}|=1$ and $\mathcal{H}_{\mathcal{B}}=1$.
@@
Note that this is a special case of the game where we have specific assumptions on the human’s preference on the team behavior and access to human’s actions. While these assumptions are quite expensive to meet, they play an important role in studying Human-AI Team PbRL algorithms. Specified Orchestration provides a loose upper-bound performance by assuming away other challenging aspects of Human-AI interaction except preference elicitation, reward learning and subsequent policy learning.
5 Team Cooperation Human Subject Study
Motivation
To formally investigate 1) the significance of Human Flexibility in Human-AI teaming, and, 2) the utility of Specified Orchestration in real-time scenarios, we conduct a user study (approved by our local Institutional Review Board (IRB)) using a modified Overcooked domain (Carroll et al. 2019), as shown in Figure 1. In particular, we compare the scenario where the human adapts to a Helper Agent (referred to as the adaptive case) and is allowed to choose from a set of strategies to the scenario of Specified Orchestration where the human has a fixed strategy that does not change during runtime (referred to as the rigid case). We refer readers to Appendix B for additional details on the study design.
Evaluation Metrics
We use a comprehensive set of metrics to evaluate both the objective task performance and the subjective user experience in our study for both cases: Task Success, Adaptation Success and Temporal Metrics such as - ”Total Time” measured the duration between the first and last action timestamps, providing insights into overall task completion and ”Active Session Time” captured the duration from web-page entry to task completion or exit, excluding instructional reading time to measure task-related engagement of the users. To evaluate frustration, temporal and cognitive load on the human in the Human-AI team setting, we take subjective feedback from the users inspired by the NASA-TLX study (Hart 2006). Here, we discuss our study findings across each of these metrics.
Study Results
Adaptation and Task Success:
Users preferred to follow their own established strategy rather than adapting to the robot’s actions. Only 6 out of 20 participants adapted by picking the bowl from the passing station. Out of these, only 2 completed the task, suggesting that adaptation does not guarantee successful task completion. Interestingly, all 10 users who did not adapt to the robot’s strategy and picked the bowl from their dispenser were able to successfully complete the task, implying that sticking to their own strategy, despite being different from the agent’s, led to task success.
Time spent:
In the adaptive scenario, users expended a total time of 1 minute and 33 seconds, a duration significantly longer than 46 seconds recorded in the rigid case. This discrepancy strongly suggests that users engaged in the adaptive setting invest a considerable amount of time in decision-making at each step, compelled by the need to select from multiple policies. Conversely, users in the rigid case adhered to a fixed policy, eliminating the need for continuous adaptation to the helper agent. The absence of this adaptability appears to significantly enhance efficiency, as evidenced by achieving task success within a shorter time frame. Therefore, we observe that human Flexibility has a strong impact on time taken for the execution of the task as following a predetermined policy streamlines the decision-making process. Examining active session times further underscores these findings (see Appendix B).
Frustration and cognitive load:
Users indicated a higher overall mental demand in the adaptive case, suggesting that adapting to the helper agent increased the cognitive workload compared to the more predetermined rigid case. Specifically, generating actions in the adaptive scenario were reported as very demanding, underscoring the challenges users faced in real-time decision-making and anticipating the actions of an adaptive agent. In assessing frustration, users reported moderate frustration levels in the adaptive case, particularly in deciding their next action, understanding the other agent’s intentions, adapting to ensure task completion and relying on the other agent. Conversely, frustration levels were notably lower across all parameters in the rigid case, indicating a smoother and less frustrating experience. In summary, while the adaptive case introduced a heightened mental demand and cognitive load, participants perceived the rigid case of Specified Orchestration as more efficient and less frustrating.
Figure 2: MA Highway domain - human and the AI agent cars are shown one behind the other (see Appendix C).
Figure 3: MA MuJoCo(L to R)- Cheetah, Ant, Walker, Swimmer, & Hopper. The action space for the joints in each of the domains has been split between the human and the AI agent (see Appendix C).
6 Experiments & Results
Figure 4: Learning curves on MA Highway-Right (row 1) and MA MuJoCo - Cheetah (row 2): comparing (from L to R) (a) Human Flexibility on multiple $\pi_{H}$, with Specified Orchestration case that assumes a single $\pi_{H}$ and complete access to it, (b) agent’s 0% access to $\pi_{H}$, and (c) partial access to $\pi_{H}$; as measured on the episodic returns. The solid lines and shaded regions represent the mean and standard deviation, respectively, across three runs.
Figure 5: Learning curves on MA MuJoCo domains (from L to R) - Walker, Hopper, Ant, Swimmer: comparing PbRL algorithms under Specified Orchestration, as measured on the episodic returns. The solid lines and shaded regions represent the mean and standard deviation, respectively, across three runs.
@@ 6 Experiments & Results
In this section we will discuss our domain suite (Section 6), followed by the experiment design setting (Section 6), and our experiments evaluating multi-agent PbRL across Human Flexibility (Section 6), Access to Human Policy, and discussing the special case of Specified Orchestration (Section 6). Further, our discussion will highlight the challenges associated with using PbRL for Human-AI team Cooperation Game and empirically answer the following questions :
@@
    Human Flexibility: What is the relative performance when the human partner can choose from multiple possible policies to follow, for accomplishing the team task?
@@
    Access to Human Policy: How do PbRL algorithms perform with a a) budgeted and b) no access to $\pi_{H}$?
@@
    Specified Orchestration: What is the performance of the PbRL algorithms with maximal assumptions on access to human policy and $\epsilon=1$-H-flexible human?
@@
We refer the readers to the Appendix (Sections A, C, D, and E) for details on algorithms, the human policy and preference oracles, domain construction, hyperparameters used for the experiments and additional results, respectively.
Evaluation Domains
MA Highway:
We design a set of three multi-agent continuous control environments on top of the Highway environment (Leurent 2018). As shown in Figure 2, we construct a three-lane highway where the human car (in orange) is on the top-most lane, and the AI agent (in purple) has to overtake and merge back into the human car’s lane by increasing its speed and avoiding collisions with any other car in the scene. Analogous to the Overcooked domain, MA Highway also does not require forced cooperation for task completion.
MA MuJoCo:
Following prior works in constructing multi-agent environments (Peng et al. 2021) in continuous control domains and PbRL setups (Lee et al. 2021), we modify the MuJoCo-based locomotion environments (Todorov, Erez, and Tassa 2012) for the human-AI teaming tasks which necessitate cooperation. We construct multi-agent settings for five environments, as shown in Figure 3.
Adapting PbRL for Human-AI Team setting
For our experiments, we assume a human policy set $|\Pi|$ from which the human agent can select a single (or multiple) policies for themselves, and the AI agent has the access to the human policy during training with a budget $\mathcal{H}_{\mathcal{B}}$. Following prior works (Strouse et al. 2021; Yu et al. 2023), when a new episode begins, we uniformly sample a human policy $\pi_{H}$ from $|\Pi|$. Note, if the human only has a single policy, i.e., in the case of Specified Orchestration, then we use the same $\pi_{H}$ for every episode. During the policy learning phase for the AI agent if the agent has remaining budget to request actions from human policy, we sample $a_{H}$ from $\pi_{H}$, or else, the AI agent assumes a random action on behalf of the human. Again, in the case of Specified Orchestration, the budget $\mathcal{H}_{\mathcal{B}}$ allows the AI agent to access human actions till the end of the training (see Algorithm 1).
Impact of Human Flexibility
To study the impact of Human Flexibility (our Q1) on PbRL performance we operate under the assumption that agent has complete access to the human policy (i.e. human executes their policy during agent’s training). We assume that for each episode of learning, the human agent can sample 1-out-of-N (3, in our case) policies to provide the actions to the AI agent. From the results shown in the first column of Figure 4, we note that the performance comparatively drops as in the case where we assumed a single $\pi_{H}$ and complete access to it. The decline in performance is amplified in the more complex case of MA MuJoCo Cheetah where cooperation is forced amongst the two agents. We conclude that the team performance is the highest in the case of Specified Orchestration, where the human agent has a single $\pi_{H}$ and the AI agent has complete access to the same.
Observed Challenges:
Impact of Access to Human Policy
No Access to Human policy:
In reference to Question (2), we report the performance of PEBBLE (Lee, Smith, and Abbeel 2021), RUNE (Liang et al. 2022), and SURF (Park et al. 2022) on the three MA Highway and MA MuJoCo - Cheetah domains in second column of Figure 4. None of the algorithms have any access to $\pi_{H}$, and thus, have to ‘imagine’, i.e., assume random actions for the human agent. In all the four domains and three algorithms, we note that the performance is significantly worse as compared to the case when the AI agent has complete access to $\pi_{H}$. Utilizing explicit modeling of human policy is a natural next step and an important future work, but it is beyond the scope of this work.
Restricting Access to Human Policy:
Furthermore, we first investigate the case where we restrict the AI agent’s budget for accessing the human policy $\pi_{H}$ to 50% for MA Highway domain, and 10% for MA MuJoCo - Cheetah (third column of Figure 4). We allocate this entire budget in the first half of the training. For example, in the case of MA MuJoCo - Cheetah domain, the AI agent has access to $\pi_{H}$ up to 100k steps, and assumes random actions sampled for the remaining training steps. Expectedly, we note that the performance plummets significantly as soon as the human agent’s actions are not available for all the domains.
Specified Orchestration - complete access to $\pi_{H}$:
Finally, to answer Question (3) supporting our discussion on Specified Orchestration, we report the performance of all three PbRL algorithms in Figures 4 and 5. We note from each of the plots, that all three algorithms show the optimal returns. From Figure 5, we also note that RUNE consistently performs better than SURF and PEBBLE in these multi-agent settings.
Observed Challenges:
From the experiments with 0% and partially restricted access to the human partner policy, we observe that reward learning is severely impacted. Since Specified Orchestration assumes maximal information on the human partner’s policy and preferences, the team performance is optimal. This highlights the potential of using Specified Orchestration as a means to obtain a loose upper bound on the Human-AI team PbRL Cooperation Game performance. Further, unless the AI agent explicitly models human policy (and preferences) with restricted access, both reward and policy learning are challenging problems.
7 Conclusion
We perform a first exploration of Preference-based Reinforcement Learning for Human-AI team setup. We introduce a Human-AI PbRL Cooperation game and argue across two critical dimensions, namely - Human Flexibility and AI agent’s Access to Human Policy. In our setup, both the agents are rewarded according to human’s reward function for the team comprising of human’s objectives for the team, sub-objectives for themselves and the robot, and human preferences on team behavior. To this end, we evaluate PbRL algorithms on our domain suite consisting of simpler Highway and more complex robot locomotion domains. From our experiments across the different cases of Human Flexibility and AI agent’s Access to Human Policy, we conclude that Specified Orchestration, while assuming the maximal amount of information on the human, shows notable performance across all domains. However, this can also be seen as one of the key limitations as it may not always be the case that the human can easily compute and provide their policy to the AI agent (or be available to act during the training procedure of the AI agent). Secondly, while this approach allows for optimal team performance of an AI agent with a specific human, it may be hard to generalize the learnt reward function (and hence, the learnt policy) to multiple human partners. Thirdly, even the case of Specified Orchestration assumes the human agent to act rationally and within the confines of pre-defined strategies, which may not always hold true in real-world scenarios where human behavior can be predictable. Lastly, as in the case of single-agent PbRL works where human is just an observer, the problem of expensive human interaction with the AI agent is carried over to our setting. To conclude, our work provides researchers with a first exploration of PbRL beyond single-agent setup, towards human-AI teaming.
Appendix A PbRL Algorithm
We present our modified PbRL algorithm for the human-AI team setting, which uses the PEBBLE algorithm (Lee, Smith, and Abbeel 2021) as a backbone. Note, we make the same modifications for experiments using RUNE (Liang et al. 2022) and SURF (Park et al. 2022) as these algorithms also use PEBBLE as their PbRL backbone. We do not require any change to the model architectures (including reward approximator and policy approximator). The updates are highlighted in red (see Algorithm 1) and can be easily incorporated into any online single agent PbRL algorithm.
Following past works in Preference based Reinforcement Learning, we use a simulated oracular human policy to provide the actions on behalf of the human agent. As suggested in (Lee et al. 2021) we use a scripted preference labeling function that provides preference labels on behalf of the human agent to the PbRL agent. We train SAC (Haarnoja et al. 2018) by assuming centralized controller for human and AI agent using underlying task reward available with the environment package for MuJoCo (Towers et al. 2023) and handcrafted reward function for Highway. We take the actions corresponding to the joint angles under control of the human as human actions and rest for the AI agent’s actions expected by the human as part of the compatible policy pair. Preference labels are generated based on hand-crafted rewards for MA Highway and environment packaged rewards (which serve as team reward) for MA MuJoCo. We refer the readers to the appendix for further details.